# pasta: Pattern Analysis for Spatial Omics Data

### v2

@@ Discussion
In this work, we presented multiple options for the exploratory data analysis of spatial omics data. We introduced the foundations of both point pattern and lattice data analysis and showed the usability of both frameworks in accompanying vignettes.
Fundamentally, point pattern and lattice data analysis are two distinct paths to analysing spatial omics data. Analysts should be aware of the assumptions and interpretations that can be drawn from each approach. In lattice data analysis, we regard the spatial coordinates as fixed and focus on the dependence between neighbours, i.e., weight matrices that reflect these neighbour interactions. In contrast, point pattern analysis concentrates on the underlying mechanisms that generate stochastic spatial patterns (i.e., clustered, regular or spaced arrangements). Therefore, the choice of the analysis framework depends not only on the data modalities, but also on the research question [anselin2024introduction, cressieStatisticsSpatialData1993, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023, daleSpatialAnalysisGuide2014]. Meanwhile, there are dualities between the data modalities from spatially-resolved measurements (c.f., Figure 1). Using cell segmentation, imaging data can be approximated as a points via their centroids. Point locations are stochastic and can therefore be regarded as a realisation of a point process and can be studied using point pattern analysis. On the other hand, we can also view segmentations as an irregular lattice (and aggregate expression per cell). Similarly, HTS-based data with subcellular resolution (e.g., Visium HD [oliveiraCharacterizationImmuneCell2024]) can likewise be interpreted as a regular lattice of (sub)-cellular locations or irregular lattice from segmentations [polanskiBin2cellReconstructsCells2024] or the reconstructed cells can then be approximated by their centroids as points and further analysed using point pattern analysis.
We have focused here on spatial statistics and highlighted two main streams of analysis. Another popular way to represent spatial arrangements is spatial graphs [aliGraphCompassSpatialMetrics2024, windhagerEndendWorkflowMultiplexed2023]. The methods described in the section on lattice data are closely connected to spatial graphs, especially in the construction of a weight matrix. Point pattern analysis concepts are also related to spatial graph via edge rules. For example, Ripley’s can be interpreted as a graph problem with a distance threshold edge rule, as explained in [daleSpatialAnalysisGuide2014, p. 136].
@@
There are various other limitations when studying spatially-resolved omics data. On the technical side, imperfect sections can lead to artifacts or non-overlapping FOVs, which can introduce missing data. Furthermore, cells are never perfectly arranged within a histological section, thus cells overlapping in the z-axis are to be expected. In terms of analysis, two-dimensional views neglect processes that occur in three dimensions. For example, in lattice data we might neglect an important contiguous neighbour in an adjacent slice. The same applies to point pattern analysis where we consider only processes in two dimensions even though the underlying biological process likely happens in three dimensions. Both lattice data and point pattern analysis extend conceptually to 3D, however with increased computational complexity. Moreover, in point pattern analysis, we often implicitly assume that patterns are invariant to rotations (see box Assumptions). This assumption is not anatomically accurate for organs and tissues with layered structures, such as the brain. There are other limitations arising from the concept of spatial metrics. Lattice data analysis methods were originally designed for rather low dimensional datasets, while spatial omics often deals with thousands of features (e.g., full transcriptome). Similarly, point pattern methods usually only allow for pairwise comparison of features, which can lead to a high number of cross comparisons e.g., when there are a dozen cell types. Moreover, with newer technologies, the size of the biological sample and resolution of measurements are constantly expanding, which can lead to computational performance issues (e.g., to store cell segmentations in memory to construct weight matrices). Since point pattern based methods only store coordinates, they usually scale better in terms of memory requirements.
@@
In our accompanying vignette, we discuss concepts and assumptions in more detail using biological examples with R and Python. For point pattern analysis, we mainly use R due to the lack of comprehensive Python libraries available for point pattern analysis. Since the R package Voyager is an extensive framework and resource for lattice data, we build on it for the lattice data analysis part of our work. In addition, we provide vignettes for point pattern analysis to complement existing discussions [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023].
Outlook
@@
The spatial metrics as described in point pattern and lattice data analysis offer interesting ways to quantify spatial patterns in molecular biology. There are still open research questions including the correct application (e.g., scale and homogeneity) of methods to a given biological context, the high dimensional nature of spatial omics data and the effect of preprocessing steps that come upstream of the application of spatial statistics. Furthermore, comparisons of spatial metrics across multiple samples and conditions and parametric spatial models offer interesting options for future work.
@@
Code and instructions to retrieve data to reproduce figures and results can be found under https://github.com/robinsonlabuzh/pasta-manuscript, code licensed under the GPLv3 terms. The vignette and source code can be accessed via https://robinsonlabuzh.github.io/pasta licensed under CC-BY-4.0 terms for the text and GPLv3 license for any code.
@@ Acknowledgments
We thank all members of the Robinsonlab for constructive feedback on the manuscript and vignette. In particular, we thank Peiying Cai, Alice Driessen, Reto Gerber, Pierre-Luc Germain, Maruša Kodermann, Vladyslav Korobeynyk, Siyuan Luo, Giulia Moro, Emanuel Sonder, Jiayi Wang and David Wissel for their careful reading and input to the vignette. The Python examples in the vignette were developed as part of the SpaceHack 2024 event. We thank the SpaceHack organisers and participants.
@@
ME: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; SG: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; HLC: Conceptualisation, Vignette, Writing - Review & Editing; IM: Methodology, Writing - Review & Editing; MK, Software (Vignette), Writing - Review & Editing; RF: Methodology, Writing - Review & Editing; MDR: Conceptualisation, Methodology, Writing - Original Draft, Supervision, Funding acquisition.
@@
Supplementary Figures

### v3

@@ Discussion
In this work, we presented multiple options for the exploratory analysis of spatial omics data. We introduced the foundations of both point pattern and lattice data analysis and showed the usability of both frameworks in accompanying vignettes.
Fundamentally, point pattern and lattice data analysis are two distinct paths to analyzing spatial omics data. Analysts should be aware of the assumptions and interpretations that can be drawn from each approach. In lattice data analysis, we regard the spatial coordinates as fixed and focus on the dependence between neighbours, i.e., weight matrices that reflect these neighbour interactions. In contrast, point pattern analysis concentrates on the underlying mechanisms that generate stochastic spatial patterns (i.e., clustered, regular or spaced arrangements). Moreover, the type of mark associated with spatial locations is another decisive factor in analysis. Lattice data analysis is most suited when the mark of interest is continuous, while point pattern analysis is the case for categorical marks. However, small nuances exist with lattice data analysis methods for categorical data and point pattern methods for continuous data. Therefore, the choice of the analysis framework depends not only on the data modalities, but also on the research question [anselin2024introduction, cressieStatisticsSpatialData1993, baddeleySpatialPointPatterns2015, pebesmaSpatialDataScience2023, daleSpatialAnalysisGuide2014]. Meanwhile, there are dualities between the data modalities from spatially-resolved measurements (c.f., Figure 1). Using cell segmentation, imaging data can be approximated as points via their centroids. Point locations are stochastic and can therefore be regarded as a realization of a point process and can be studied using point pattern analysis. On the other hand, we can also view segmentation as an irregular lattice (and aggregate expression per cell). Similarly, HTS-based data with subcellular resolution (e.g., Visium HD [oliveiraCharacterizationImmuneCell2024]) can likewise be interpreted as a regular lattice of (sub)-cellular locations or irregular lattice from segmentation [polanskiBin2cellReconstructsCells2024] or the reconstructed cells can then be approximated by their centroids as points and further analyzed using point pattern analysis.
We have focused here on spatial statistics and highlighted two main streams of analysis. Another popular way to represent spatial arrangements is as spatial graphs [aliGraphCompassSpatialMetrics2024, windhagerEndendWorkflowMultiplexed2023]. The methods described in the section on lattice data are closely connected to spatial graphs, especially in the construction of a weight matrix. Point pattern analysis concepts are also related to spatial graphs via edge rules. For example, Ripley’s can be interpreted as a graph problem with a distance threshold edge rule [daleSpatialAnalysisGuide2014, p. 136].
Finding the correct scale of analysis
@@
There are various other limitations when studying spatially-resolved omics data. On the technical side, imperfect sections can lead to artifacts or non-overlapping FOVs, which can introduce missing data. Furthermore, cells are never a perfectly monolayer within a histological section, thus cells overlapping in the -axis are to be expected. In terms of analysis, two-dimensional views neglect processes that occur in three dimensions. For example, in lattice data we might neglect an important contiguous neighbour in an adjacent slice. The same applies to point pattern analysis where we consider only processes in two dimensions even though the underlying biological process likely happens in three dimensions. Both lattice data and point pattern analysis extend conceptually to 3D, however with increased computational complexity. Moreover, in point pattern analysis, we often implicitly assume that patterns are invariant to rotations (see Box Importance of scale and uniformity in point pattern analysis). This assumption is not anatomically accurate for organs and tissues with layered structures, such as the brain. Moreover, there are other limitations arising from the concept of spatial metrics. Lattice data analysis methods were originally designed for rather low dimensional datasets, while spatial omics often deals with thousands of features (e.g., untargeted transcriptome). In addition, the interpretation of a spatial correlation metric depends on the measure of similarity that is compared in a local neighbourhood. In a multivariate setting, a straightforward interpretation of this similarity measure is often not possible. Similarly, point pattern methods usually only allow for pairwise comparison of features, which can lead to a high number of cross comparisons. Moreover, with newer technologies, the size of the biological sample and resolution of measurements are constantly expanding, which can be computationally expensive. Since point pattern based methods only store coordinates, they usually scale better in terms of memory requirements.
@@
In our accompanying vignette, we discuss concepts and assumptions in more detail using biological examples. For point pattern analysis, we mainly use R due to the lack of comprehensive Python libraries available for point pattern analysis. Since the R package Voyager is an extensive framework and resource for lattice data, we build on it for the lattice data analysis part of our work. In addition, we provide vignettes for point pattern analysis to complement existing discussions [zormpasMappingTranscriptomeRealizing2023, mosesVoyagerExploratorySinglecell2023].
Future opportunities
@@
The spatial metrics as described in point pattern and lattice data analysis offer means to quantify spatial patterns in molecular biology. There are still open research questions including the correct application (e.g., scale and homogeneity) of methods to a given biological context, the high dimensional nature of spatial omics data and the effect of preprocessing steps that come upstream of the application of spatial statistics. Furthermore, comparisons of spatial metrics across multiple samples and conditions and parametric spatial models offer interesting options for future work.
Methods
All analysis were conducted using R version 4.5.0 and Bioconductor version 3.21 [gentleman2004bioconductor]. Results were visualized using ggplot2 [wickhamGgplot2ElegantGraphics2016].
[…]
The mouse brain datasets shown in Figure 1 (Technology) were taken from Shi et al. (STARMAP Plus) [shiSpatialAtlasMouse2023] [shiSpatialAtlasMouse2023] and 10x (Visium mouse) [10xVisium_Mouse]. Clustering was performed using Banksy [singhalBANKSYUnifiesCell2024].
Lattice data analysis was performed using the package Voyager [mosesVoyagerExploratorySinglecell2023] and spdep [pebesmaAppliedSpatialDataAnalysis2013], point pattern analysis was conducted using the packages spatialFDA [emonsSpatialFDAToolSpatial2025] and spatstat [baddeleySpatstatPackageAnalyzing2005]. Biologically relevant observation windows were determined using the package sosta [gunzSostaPackageAnalysis2025].
@@ Methods
Code and instructions to retrieve data to reproduce figures and other results can be found under https://github.com/robinsonlabuzh/pasta-manuscript, code licensed under the GPLv3 terms. The vignette and source code can be accessed via https://robinsonlabuzh.github.io/pasta licensed under CC-BY-4.0 terms for the text and GPLv3 license for any code.
Versioned source code and rendered vignettes available under https://doi.org/10.5281/zenodo.15488040.
@@ Acknowledgments
We thank all members of the Robinsonlab for constructive feedback on the manuscript and vignette. In particular, we thank Peiying Cai, Alice Driessen, Reto Gerber, Pierre-Luc Germain, Maruša Kodermann, Vladyslav Korobeynyk, Siyuan Luo, Giulia Moro, Emanuel Sonder, Jiayi Wang and David Wissel for their careful reading and input to the vignette.The Python examples in the vignette were developed as part of the SpaceHack 2024 event. We thank the SpaceHack organisers and participants. Furthermore, we would like to thank Daria Lazic for helpful comments on the manuscript.
@@
ME: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; SG: Conceptualisation, Methodology, Software (Vignette), Analysis, Visualisation, Writing - Original Draft; HLC: Conceptualisation, Vignette, Writing - Review & Editing; IM: Methodology, Software (Vignette), Writing - Review & Editing; MK, Software (Vignette), Writing - Review & Editing; RF: Methodology, Writing - Review & Editing; MDR: Conceptualisation, Methodology, Writing - Original Draft, Supervision, Funding acquisition.
@@
Supplementary data
Technology | Main feature of interest | Lattice data | Point pattern | |||
Visium [stahlVisualizationAnalysisGene2016] | transcriptomics | spots (regular) | not applicable | |||
Slide-seq V1& V2 [rodriquesSlideseqScalableTechnology2019, stickelsHighlySensitiveSpatial2021] | transcriptomics |
segmented cells | |||
Visium HD [oliveiraCharacterizationImmuneCell2024] | transcriptomics |
segmented cells | |||
Stereo-seq [chenSpatiallyResolvedHighly2015] | transcriptomics |
segmented cells | |||
Xenium [janesickHighResolutionMapping2023] | transcriptomics | segmented cells (irregular) |
CosMx [chenSpatiallyResolvedHighly2015] | transcriptomics | segmented cells (irregular) |
MERFISH [chenSpatiallyResolvedHighly2015] | transcriptomics | segmented cells (irregular) |
IMC [giesenHighlyMultiplexedImaging2014] | proteomics |
segmented cells | |||
MIBI-TOF [kerenMIBITOFMultiplexedImaging2019] | proteomics |
segmented cells | |||
CODEX [blackCODEXMultiplexedTissue2021] | proteomics |
4i [gutMultiplexedProteinMaps2018] | proteomics |
![[Uncaptioned image]](x9.png)
(Figure legend on next page.)